OpenStudy (ash02):

Find the slope: (3,-5) (-3,5) Slope=

OpenStudy (ash02):

I got 5/-3 is it right? Sorry I'm behide a year:(

OpenStudy (markkboyy):

m=−5/3

OpenStudy (ash02):

How?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) so, m = (5-(-5))/(-3-3) = 10/-6 = -5/3 yes you're correct :)

OpenStudy (ash02):

-5/3 I got 5/-3

OpenStudy (ash02):

Oh Nvm thanks

OpenStudy (ash02):

Lol sorry

hartnn (hartnn):

they're same :) welcome ^_^

OpenStudy (ash02):

How about this one? Find the slope: (-1,-5) (-2,5) Slope=

OpenStudy (ash02):

I got 10/1

hartnn (hartnn):

(5-(-5))/ (-2-(-1)) = 10/(-2+1) = 10/-1 = -10 check your denominator?

OpenStudy (ash02):

-10 okay thanks